News Alert: AS OF 1400 HOURS 18 JUNE 2013 ACTIVATION LEVEL: CRISIS ACTION TEAM AS OF 1500 HOURS 14 JUNE 2013
Posted: 6/18/2013 3:14:50 PM
Louisiana Business Emergency Operations Center




 



022-13 Pipeline Leak Franklin Parish (Delhi) 13 June 2013



 



UPDATED INFORMATION IS IN BOLD
RED



 



SITUATION



At 1402 hours 13 June 2013, notification was received of a
pipeline leak in Franklin Parish near the community of Delhi. Denbury Resources
and Trans Canada are the two companies that have pipelines in the area. Initial responsible party was thought to be Trans Canada, but
has been recently confirmed that this pipeline belongs to Denbury.
 



 



18 June 2013



The four wells around the release
locations were pumped with mud through the night with little to no decrease in
the uncontrolled flow.  The release began in one location and has now
migrated under a pipeline right of way.  There were five boils in the right
of way yesterday evening and now there are twelve. The right of way is home to
two (2) 30 inch and one (1) 36 inch compressed natural gas pipelines. 
These pipelines were installed in the late 1950's and the early 1960’s. These
pipelines operate at 750 psi and approximately 20 million cubic feet per hour
pass this possibly compromised area every hour. A decision has been made by
LSP/ESU at Denbury’s suggestion to block in and blow down these three pipelines
until their integrity can been better assessed.  Plans are underway to
mitigate this blow down. There will be highway and railway closures during this
period of time. At approximately 1030 hours, an entry was made by Denbury,
USES, and Boots and Coots personnel.  Monitoring data identified the LEL
at 60%, the O2 readings were 14.6%, and the CO2 readings were maximized the
meter at 50,000 ppm.  Further entries will be made after the compressed
gas pipelines have been depressurized. Updates will come when future entry
teams obtain information.



 



17 June 2013



As of 0943 hours, the incident was ongoing. Boots and Coots
had made it to within 60 yards of the source. Air monitoring is ongoing with no
increase in readings. Pumping operations to bridge the source is ongoing with
progress being made.



 



16 June 2013:



At 0507 hours it was reported that workers were backed up
from the collection site due to a raised VOC and CO2 reading from air monitors.
Collection will continue once reading drop again. Boots and Coots continue to
pump heavy water and mud into the system to try to stop the leak. As of 1046
hours, the readings were still too high to approach the well site.



 



15 June 2013



As of 0822 hours, it was reported that the incident action
plan for 15 June was finished. Continuous air monitoring was ongoing. Readings
taken at all evacuated areas were good. The evacuation is still in effect but
was being evaluated. The road closure for Antely Road was still in place. The
evacuation order was lifted at 1400 hours with each evacuee to be escorted to
their homes by a Sheriff’s Deputy and a representative from CTEH to conduct air
quality testing at each residence to ensure the residence was safe. Boots and
Coots began dumping heavy water into the system to try and shut the well in.
Aerial surveillance showed the source was still leaking what appeared to be
drilling mud which was now flowing and not cascading as it was on 14 June.



 



As of 2000 hours, a containment dam was installed.



 



14 June 2013



As of 1634 hours 14 June, it was reported that Denbury is
actually the responsible party. A well was blown out and Boots and Coots were
engaged to plug the well with the work planned for 15 June. Evacuations were
being reevaluated.  As of 1857 hours, the incident was ongoing with
evacuations still in place. Air monitoring was ongoing by USES and CTEH.
Indications were that this incident would continue for several days. LSP, DEQ,
Richland Parish Sheriff’s Office and both Richland and Franklin OEP directors
were on site.



 



As of 0800 hours 14 June, the wrong line was blown the night
of 13 June and the leak is ongoing and pressure readings on the line are not
dropping. Trans Canada has a contractor en route. Antely Road remains closed
and the evacuation is still in effect.



 



13 June 2013



As of 1552 hours 13 June, LSP Hazmat reported that monitoring
and identification of the leaking line was ongoing and La Hwy 132 remained
closed as a precautionary measure.



 



There have been no reports of any explosions or fires
associated with the leak.



 



There have been no injuries or fatalities reported for this
incident.



 



Initially, 7 homes were evacuated as a precautionary measure
and one road (Antley Road in Franklin Parish) was closed. As of 1630 hours, 5
additional homes were evacuated. Residents are being set up in a local hotel by
Trans Canada. The evacuation order was lifted at 1400 hours 15 June with
residents escorted back home and their residences tested for air quality by
CTEH.



 



No requests for state assistance have been submitted.



 



ACTIONS BY PARISHES



OEP Directors for both Franklin and Richland Parish are
engaged.



 



ACTIONS BY STATE AND OTHER AGENCIES



GOHSEP



GOHSEP Region 8 Coordinator is assisting the parish affected
and providing reports as needed.



 



WebEOC incident 022-13 Pipeline Leak Franklin Parish (Delhi)
13 June 2013 has been created to support this event.



 



GOHSEP Crisis Action Team was activated at 1500 hours 14
June 2013 and will remain activated until further notice.



 



LSP



Personnel responded to the incident site to monitor and
report as needed and will maintain personnel on site until secured.



 



Louisiana State
Police is currently conducting aerial reconnaissance to assist LSP Emergency
Services personnel on scene with the assessment of the ongoing pipeline gas
leak in Franklin Parish
.



 



LDEQ



Personnel are performing air monitoring and will continue to
do so until the site is secure



 



DOTD (Dist. 58)



Personnel provided two (2) DMSs
and four (4) Type III Barricades at the request of the Franklin Parish
Sheriff’s Office (FPSO) on June 13, 2013.  DMSs were installed at the
intersection of LA 132/857 and LA 132/17.  Barricades were placed at
physical closure manned by the FPSO.  Road closure information was
provided to TMC by the District and removed along with traffic control devices
on June 15, 2013.



 



GOHSEP Crisis Action Team will monitor, report and assist
with this incident as needed.



 



 


The LA BEOC Portal is designed and developed by the NIMSAT Institute under the Cooperative Endeavor Agreement with the Louisiana Governor's Office of Homeland Security and Emergency Preparedness (GOHSEP)

GOHSEP LED NIMSAT

© Copyright 2010-2018, NIMSAT. All rights reserved. Terms of Service. Portal Guidelines.